OpenStudy (anonymous):

A swimming pool has to be drained for maintenance. The pool is shaped like a cylinder with a diameter of 8 m and a depth of 2.1 m . If the water is pumped out of the pool at the rate of 11 m^3 per hour, how many hours does it take to empty the pool? Use the value 3.14 for π , and round your answer to the nearest hour.

OpenStudy (danjs):

volume = 3.14 *4*4*2.1

OpenStudy (danjs):

time = volume /rate

OpenStudy (danjs):

105.5/11 about 9.6 hrs

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

A street that is 179m long is covered in snow. City workers are using a snowplow to clear the street. The snowplow has tires that are1.3m in diameter. How many times does a tire have to turn in traveling the length of the street? Use the value 3.14 for π . Round your answer to the nearest tenth. Do not round any intermediate steps. Is this the same as the other problem? @DanJS

OpenStudy (danjs):

hi

OpenStudy (danjs):

So this is asking you how many times a tire with diameter of 1.3m turn around to cover 179m distance.

OpenStudy (danjs):

the circumference of a tire is : C = 2*3.14*radius

OpenStudy (danjs):

how many of those circumferences are in the total distance to travel of 179 meters?

OpenStudy (danjs):

i used 1.8 for diameter, it is 1.3...1 sec

OpenStudy (anonymous):

what would be the formula for the whole thing?

OpenStudy (danjs):

oops

OpenStudy (danjs):

\[\frac{ distance }{ circumference } = \frac{ 179 }{ 2*\pi*\frac{ 1.3 }{ 2 } }\]

OpenStudy (danjs):

=179/ 4.082 Total rotations for a tire with a diameter of 1.3m to travel a distance of 179m is -- about 43.85